OpenStudy (anonymous):

Determine the sum of a geometric series whose first term is 12, whose common ratio is -1/4 and whose number of terms is infinite.

OpenStudy (anonymous):

i tried what u said last time but got a answer that was not on there wanted to make sure it was none of they above

OpenStudy (anonymous):

the*

OpenStudy (anonymous):

which one , the previous?

OpenStudy (anonymous):

yeah i tried what u said on the last one

OpenStudy (anonymous):

\[\frac{c}{1-r}\]

OpenStudy (anonymous):

is it 16?

OpenStudy (anonymous):

\[\frac{12}{1+\frac{1}{4}}\]

OpenStudy (anonymous):

i got 12.25? thst right?

OpenStudy (anonymous):

but i did 12/1-(-1/4)

OpenStudy (anonymous):

\[\frac{12}{\frac{5}{4}}\] \[12\times \frac{4}{5}=9.6\]

OpenStudy (anonymous):

oh yeah 1+ 1/4 is not 3/4

OpenStudy (anonymous):

did i mess up?